# VidLaDA: Bidirectional Diffusion Large Language Models for Efficient Video Understanding

### v1

@@ E.3 CoT Inference Visualizations
E.4 Ablation Studies Details
In this section, we provide a comprehensive analysis of the proposed MARS-Cache. We validate our design choices regarding anchor token searching, modality-wise, and layer-wise refresh rates. Unless otherwise stated, all ablation experiments are conducted on EgoSchema (subset) and MLVU (test) to evaluate reasoning and long-context capabilities, respectively. The model layers are simply divided into four groups: Group 1 (layers 0-7), Group 2 (8-15), Group 3 (16-23), and Group 4 (24-31). The detailed settings for the ablation study are provided in Appendix F.5.
E.4.1 Effectiveness of Anchor Token Searching
@@
E.4.2 Modality-Wise Refreshing Strategy
@@
E.4.3 Layer-wise Refreshing Strategy
@@
32 32 16 8 67.0 6.7 45.9 6.8
($\times 2$) 66.8 9.2 46.9 9.3
32 16 8 4 66.6 4.6 44.8 4.7
($\times 2$) 66.4 6.8 44.3 7.0
64 32 16 8 68.0 7.0 45.3 7.0
($\times 2$) 65.8 9.4 45.3 9.5
E.4.4 Overhead of Anchor Token Searching
@@
E.5 Causal Inference Visualizations
@@
Appendix F Experimental Details
F.1 Experimental Setup
Implementation Details. VidLaDA is implemented using the LLaDA-8B (Nie et al., 2025) as the LLM backbone, coupled with SigLIP2-SO400M (Tschannen et al., 2025) as the ViT, following the LLaDA-V (You et al., 2025). The model is trained following the three-stage curriculum strategy described in Section D.1. The specific training settings are listed in Table 9. Training is conducted on 32 NVIDIA H200 GPUs with a global batch size of 64. The learning rate can refer to Table 3.
@@ F.1 Experimental Setup
Datasets and Evaluation Benchmarks. We conduct a comprehensive evaluation across eight diverse benchmarks to assess VidLaDA’s capabilities in video understanding. To rigorously test long-term temporal dependencies and understanding, we employ Video-MME (Fu et al., 2025) (w/o subtitles), MLVU (Zhou et al., 2025) (reporting both Dev and Test splits), LVBench (Wang et al., 2025b), and LongVideoBench (Wu et al., 2024). These datasets challenge models with extended video contexts ranging from minutes to hours, requiring sustained attention over long sequences. For fine-grained temporal perception, we utilize MVBench (Li et al., 2024b), which comprises 20 distinct tasks requiring precise action and attribute recognition. Furthermore, we evaluate high-level reasoning and expert domain knowledge using EgoSchema (Mangalam et al., 2023), a benchmark for complex egocentric video understanding, and Video-MMMU (Hu et al., 2025), which tests multi-disciplinary professional understanding.
F.2 Intra-Frame Understanding on MME
@@
We employ an information-guided shuffling protocol to test the sensitivity of the model to the position of semantic features. First, we compute the $\ell_{2}$-norm of the output features from the vision encoder for every patch, serving as a proxy for information density. We then identify the top-$k$ tokens with the highest norms, setting $k=256$, to isolate the most semantically salient regions of the image. To manipulate the sequence order, we define a position ratio $r\in[0,1]$. The input sequence is constructed such that the selected high-norm tokens are placed starting at index $\lfloor r\times(|V|-k)\rfloor$, where $|V|$ is the total number of visual tokens. Consequently, $r=0$ positions the high-information tokens at the visual start, while $r=1$ pushes them to the visual end, with the remaining low-norm tokens filling the rest of the sequence. By evaluating the performance variance across different $r$ values, we can determine whether a model exhibits position-dependent early receptive field biases.
F.3 Inter-Frame Interaction on RexTime
@@
F.4 CoT Inference Settings
This section details the inference hyperparameters employed for the reasoning experiments using the Chain-of-Thought (CoT) framework described in Appendix D.2. Regarding visual input processing, we uniformly sample up to 32 frames from the input video.
For the intermediate reasoning analysis, we allocate a generation budget of 1024 tokens for LongVideoBench and the EgoSchema (subset), and 2048 tokens for MLVU (test), to accommodate detailed thought processes, and we report the best performance achieved. Furthermore, capitalizing on the non-autoregressive nature of the underlying Diffusion Language Model, we enable parallel decoding (Wu et al., 2025) to accelerate the inference of these extended reasoning chains. Additionally, for the MLVU benchmark, we also employ CreditDecoding (Wang et al., 2025a) to ensure decoding stability.
@@ F.4 CoT Inference Settings
F.5 Ablation Settings
Distinct from the multi-stage Chain-of-Thought framework described in Appendix F.4, our ablation studies prioritize experimental control and throughput stability to rigorously evaluate the efficiency of the proposed MARS-Cache. To decouple the efficiency metrics from the variance introduced by dynamic task routing and variable-length generation, we adopt an explicit thought-tagging strategy inspired by LLaVA-CoT (Xu et al., 2025).

### v2

@@ F.3 CoT Inference Visualizations
F.4 Architecture Comparison under Matched Fine-tuning
To disentangle architectural effects from training-related factors, we conduct a controlled fine-tuning experiment. Since re-training LLM and MLLM from scratch on the large-scale corpus and the visual instruction dataset is computationally prohibitive (Cheng et al., 2025), we compare two publicly available MLLM baselines with similar starting performance, i.e., LLaVA-OneVision as the AR baseline and LLaDA-V as the DLM baseline. We fine-tune both models using the same 10K samples and identical hyperparameters. Specifically, we freeze the ViT encoder and set the learning rate to 1e-5 for both the LLM and the MLP projector, other training settings follow Table 11. This protocol reduces differences stemming from data and optimization recipes, making the observed deltas more attributable to architectural choices.
As shown in Table 7, the AR baseline exhibits a trade-off: it improves on LongVideoBench but degrades on LVBench and MVBench. In contrast, the DLM baseline achieves consistent improvements across all benchmarks, with a particularly notable gain on Video-MME. Under matched fine-tuning conditions, these results suggest that the DLM architecture may utilize the available spatiotemporal evidence more effectively than the AR counterpart, which is consistent with our understanding-efficiency analysis in Proposition 3.2.
Table 7: Results of Architectural Comparisons. Effect of extending training for AR vs. DLM baselines under the same training dataset and identical hyperparameters.
Model #F LongVideoBench LVBench MVBench Video-MME Avg
AR Baselines
LLaVA-OneVision 32 56.5 40.7 56.7 58.5 53.1
+SFT 32 57.4 (+0.9) 38.8 (-1.9) 54.1 (-2.6) 58.6 (+0.1) 52.2 (-0.9)
DLM Baselines
LLaDA-V 32 58.2 36.4 53.1 56.4 51.0
+SFT 32 58.9 (+0.7) 36.6 (+0.2) 53.7 (+0.6) 58.5 (+2.1) 51.9 (+0.9)
F.5 Detailed Ablation Studies of MARS-Cache
In this section, we provide a comprehensive analysis of the proposed MARS-Cache. We validate our design choices regarding anchor token searching, modality-wise, and layer-wise refresh rates. Unless otherwise stated, all ablation experiments are conducted on EgoSchema (subset) and MLVU (test) to evaluate reasoning and long-context capabilities, respectively. The model layers are simply divided into four groups: Group 1 (layers 0-7), Group 2 (8-15), Group 3 (16-23), and Group 4 (24-31). The detailed settings for the ablation study are provided in Appendix G.5.
F.5.1 Effectiveness of Anchor Token Searching
@@
F.5.2 Modality-Wise Refreshing Strategy
@@
F.5.3 Layer-wise Refreshing Strategy
@@
32 32 16 8 67.0 12.0 45.9 12.2
($\times 2$) 66.8 16.4 46.9 16.7
32 16 8 4 66.6 8.2 44.8 8.3
($\times 2$) 66.4 12.2 44.3 12.4
64 32 16 8 68.0 12.5 45.3 12.6
($\times 2$) 65.8 16.9 45.3 17.0
F.5.4 Overhead of Anchor Token Searching
@@
F.6 Causal Inference Visualizations
@@
Appendix G Experimental Details
G.1 Experimental Setup
Implementation Details. VidLaDA is implemented using the LLaDA-8B (Nie et al., 2025) as the LLM backbone, coupled with SigLIP2-SO400M (Tschannen et al., 2025) as the ViT, following the LLaDA-V (You et al., 2025). The model is trained following the three-stage curriculum strategy described in Appendix E.1. The specific training settings are listed in Table 11. Training is conducted on NVIDIA H200 GPUs with a global batch size of 64. The learning rate can refer to Table 4.
@@ G.1 Experimental Setup
Datasets and Evaluation Benchmarks. We conduct a comprehensive evaluation across eight diverse benchmarks to assess VidLaDA’s general capabilities in video understanding. First, to evaluate holistic understanding across varied contexts and durations, we employ Video-MME (Fu et al., 2025) (w/o subtitles), MLVU (Zhou et al., 2025) (reporting both Dev and Test splits), LVBench (Wang et al., 2025c), and LongVideoBench (Wu et al., 2024). These benchmarks challenge models with complex visual information and sustained contexts, assessing the model’s robustness in handling diverse video content ranging from short clips to extended sequences. Second, for fine-grained temporal perception, we utilize MVBench (Li et al., 2024b), which comprises 20 distinct tasks requiring precise action and attribute recognition. Third, we evaluate high-level reasoning and expert knowledge acquisition using EgoSchema (Mangalam et al., 2023), a benchmark for complex intent understanding, and Video-MMMU (Hu et al., 2025), which tests multi-disciplinary professional understanding.
G.2 Intra-Frame Understanding on MME
@@
We employ an information-guided shuffling protocol to test the sensitivity of the model to the position of semantic features. First, we compute the $\ell_{2}$-norm of the output features from the vision encoder for every patch, serving as a proxy for information density. We then identify the top-$k$ tokens with the highest norms, setting $k=256$, to isolate the most semantically salient regions of the image. To manipulate the sequence order, we define a position ratio $r\in[0,1]$. The input sequence is constructed such that the selected high-norm tokens are placed starting at index $\lfloor r\times|{{\bm{E}}^{\mathcal{V}}}|\rfloor$, where $|{{\bm{E}}^{\mathcal{V}}}|$ is the total number of visual tokens. Consequently, $r=0$ positions the high-information tokens at the visual start, while $r=1$ pushes them to the visual end, with the remaining low-norm tokens filling the rest of the sequence. By evaluating the performance variance across different $r$ values, we can determine whether a model exhibits position-dependent early receptive field biases.
G.3 Inter-Frame Interaction on RexTime
@@
G.4 CoT Inference Settings
This section details the inference hyperparameters employed for the reasoning experiments using the Chain-of-Thought (CoT) prompts described in Appendix E.2. Regarding visual input processing, we uniformly sample up to 32 frames from the input video. For MARS-Cache settings, we adopt the optimal hyperparameters identified in the ablation studies (Section 5.3).
For the intermediate reasoning analysis, we allocate a generation budget of 1024 tokens for LongVideoBench and the EgoSchema (subset), and 2048 tokens for MLVU (test), to accommodate detailed thought processes, and we report the best performance achieved. Furthermore, capitalizing on the non-autoregressive nature of the underlying Diffusion Language Model, we enable parallel decoding (Wu et al., 2025) to accelerate the inference of these extended reasoning chains. Additionally, for the MLVU benchmark and LLaDA-V, we also employ CreditDecoding (Wang et al., 2025a) to ensure decoding stability.
@@ G.4 CoT Inference Settings
G.5 Ablation Settings
Distinct from the multi-stage Chain-of-Thought framework described in Appendix G.4, our ablation studies prioritize experimental control and throughput stability to rigorously evaluate the efficiency of the proposed MARS-Cache. To decouple the efficiency metrics from the variance introduced by dynamic task routing and variable-length generation, we adopt an explicit thought-tagging strategy inspired by LLaVA-CoT (Xu et al., 2025).